Lancaster Lumber buys $8 million of materials (net of discounts) on terms of 3/5, net 75, and it currently pays on the 5th day and takes discounts. Lancaster plans to expand, which will require additional financing. Assume 365 days in year for your calculations.

If Lancaster decides to forgo discounts, how much additional credit could it obtain? Write out your answer completely. For example, 5 million should be entered as 5,000,000. Do not round intermediate calculations. Round your answer to the nearest cent.
$  

What would be the nominal cost of that credit? Do not round intermediate calculations. Round your answer to two decimal places.
  %

What would be the effective cost of that credit? Do not round intermediate calculations. Round your answer to two decimal places.
  %

If the company could get the funds from a bank at a rate of 7%, interest paid monthly, based on a 365-day year, what would be the effective cost of the bank loan? Do not round intermediate calculations. Round your answer to two decimal places.
  %
Answer #1

Additional credit period = (75 - 5) = 70 days

Additional credit = Raw material purchase / 365 x Additional credit period = $ 8,000,000 / 365 x 70 = $ 1,534,246.58

NominalCost = Discount 1 - Discount 365 * 100 Additional Daysof Credit

Nominal cost = 3% / (1 - 3%) x 365 / 70 x 100 = 16.13%

365 EffectiveCost = [(1+ Discount pan .) additio 1- Discount y co Credit] – 1

Effective cost = [(1 + 3% / (1 - 3%)]365 / 70 - 1 = 17.21%

EAR of a loan

EAR = (1+)-1

EAR = (1 + 7%/12)12 - 1 = 7.23%

------------------------------------